OpenStudy (anonymous):

determine the equation of a line perpendicular to this equation, at its y-intercept.

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 }e ^{x+1}\]

OpenStudy (mertsj):

Did you find the y intercept?

OpenStudy (anonymous):

yes, i got e/2

OpenStudy (anonymous):

so my point is (0, e/2). My derivative is \[\frac{ 1 }{ 2 }xe ^{x+1}\]. Would I sub in 0 for the equation to solve for the slope? because I get 0 as a slope and the back has 2/e

OpenStudy (mertsj):

You want the slope at the y intercept which is e/2

OpenStudy (anonymous):

yes, so would I sub that in instead

OpenStudy (mertsj):

Oh. That's right...you need the slope at the x value 0.

OpenStudy (anonymous):

but the slope at x value at 0 is 0

OpenStudy (mertsj):

The first derivative is the same as the original function.

OpenStudy (mertsj):

So if you replace x with 0 you do get e/2 for the slope.

OpenStudy (anonymous):

no, my first derivative is \frac{ 1 }{ 2 }xe ^{x+1}

OpenStudy (mertsj):

no

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 }xe ^{x+1} \]

OpenStudy (anonymous):

OMG I didn't take the derivative of x

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 }e ^{x+1}\]

OpenStudy (mertsj):

that's it. and if you replace x with 0 you get e/2 for the slope

OpenStudy (anonymous):

yes, ugh so stupid. Can you help with another question

OpenStudy (mertsj):

Maybe...I don't know a lot.

OpenStudy (anonymous):

Radium decays at a rate that is proportional to its mass, and has a half-life of 1590 years. If 20g of radium is present initially, how long will it take for it to decay to 90% of its its mass.

OpenStudy (anonymous):

I tried to do it, but I got a different answer

OpenStudy (mertsj):

Remind me of the exponential decay formula.

OpenStudy (anonymous):

Mf=Mi(1/2)t/h

OpenStudy (anonymous):

^t/h

OpenStudy (mertsj):

\[18=20(\frac{1}{2})^{]\frac{t}{1590}}\]

OpenStudy (mertsj):

Is that what you did?

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

\[\ln .9=\frac{t}{1590}\ln .5\]

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

\[t=\frac{\ln .9}{\ln .5}\times 1590=241.7\]

OpenStudy (mertsj):

Is that what you got?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but its not the answer my book gives

OpenStudy (mertsj):

And what is the real answer?

OpenStudy (anonymous):

5282 years

OpenStudy (mertsj):

I didn't remember that formula. Let me research the formula.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Still need help?

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

\[A=A _{0}e ^{kt}\]

OpenStudy (mertsj):

Use that to find the constant.

OpenStudy (mertsj):

\[50=100e ^{k(1590)}\]

OpenStudy (mertsj):

\[\ln .5=1590k\]

OpenStudy (mertsj):

\[k=-.0004359\]

OpenStudy (anonymous):

could you explain why I need the constant, because its the idea behind the question I don't get

OpenStudy (mertsj):

\[18=20e ^{-.0004359t}\]

OpenStudy (mertsj):

Because that is the exponential decay equation. They gave you the half-life so you could find the constant of variation. Remember the problem said it decays at a rate proportional to its mass.

OpenStudy (anonymous):

so how will I know when to do it for another question

OpenStudy (mertsj):

I guess when you don't know the constant.

OpenStudy (anonymous):

stupid question, but how will I know what the constant is?

OpenStudy (anonymous):

and if it gives me one, I know how to find it now, but I don't get how to know if i'm given one or not

OpenStudy (mertsj):

If you are given the constant, it will be identified as such.

OpenStudy (anonymous):

ok. umm could you help me with a vector question

OpenStudy (mertsj):

You might want to post it as a new question. I am not good with vectors.

OpenStudy (anonymous):

i did, but no one answered

OpenStudy (anonymous):

could I ask you a question unrelated to math

OpenStudy (mertsj):

yes

OpenStudy (mertsj):

Where is the vector question. I can look at it and see if I know how to do it.

OpenStudy (anonymous):

I have exams in a week or so. I have so much anxiety right now, that I shake when studying. I know everything, but I fear that on my exam there will be a question that I will know what to do, but I won't be able to do it. Because this has happened before and I spend the whole class working on one question and left 4 other question blank. I don't know what to do.

OpenStudy (mertsj):

Go through the test and do all the problems you know how to do. The ones you are not sure of..go back to at the end starting with the one you are most familiar with.

OpenStudy (anonymous):

what about until then, bc i'm freaking out about studying

OpenStudy (mertsj):

Do you have a study guide?

OpenStudy (anonymous):

no, just my notes, and my teacher keeps saying the exam is a killer and that if ppl study the will know what to do, but its a matter of knowing how to do it. Like I could know how to take the derivative of a ln function, but I don't know were to start.

OpenStudy (anonymous):

its just not sitting well with me, since I study for his tests and then mess up

OpenStudy (mertsj):

Why don't you repost that other question. It is closed. If you keep reposting it, eventually someone will answer it. I don't know how to do it.

OpenStudy (anonymous):

its ok, i will ask my teacher

OpenStudy (mertsj):

Are you trying to memorize how to do the problems? In math, you have to know what to memorize and what to reason out.

OpenStudy (anonymous):

no, like I know the idea behind the equations, chain rule, product rule etc. and I know how to apply them. I memorize the log properties, but not the questions.

OpenStudy (mertsj):

You will be fine. Just relax and do your best...that's all anyone can expect.

OpenStudy (anonymous):

but my best is in the 70s which is a B- to a B+. I want to get at least a A- for my university to see.